OpenStudy (anonymous):

x^2+10x+36 = (x+a)^2 +b What is the value of a and b?

Parth (parthkohli):

Simply complete the square in this case.

Parth (parthkohli):

@ash2326 don't stop :)

OpenStudy (ash2326):

@InspyR4D did you get the hint @ParthKohli gave?

OpenStudy (anonymous):

yup, thanks. I've completed the square but I don't know how to individually work out x, a or b

Parth (parthkohli):

I wouldn't like to intimidate, but look at the following formula:\[ax^2 + bx + c \implies ax^2 + bx - \left(b\over 2 \right)^2 + \left( b \over 2\right)^2 + c\]

Parth (parthkohli):

It isn't too hard when you are working with it!\[x^2 + 10x + 36 \implies x^2 + 10x - \left(10 \over 2 \right)^2 + \left( 10 \over 2\right)^2 + c \]Or,\[x^2 + 10x - 5^2 + 5^2 + c\]

Parth (parthkohli):

After that, we are left with the following:\[x^2 + 10x - 5^2 + 5^2 + 36\]We can simplify this by using the method stated below.\[x^2 + 10x + 25 - 25 + 36\]\[\implies x^2 + 10x + 25 + 11\qquad(-25 + 36 = +11)\]\[\implies (x + 5)^2 + 11\]